FIND THE MEAN OF THE SET OF DATA BELOW AND CONSTRUCT A SAMPLING DISTRIBUTION

BY SELECTING 3 SAMPLES AT A TIME: 7 10 14 17 20


1
Expert's answer
2022-05-13T02:57:35-0400

We have population values 7,10,14,17,20, population size N=5 and sample size n=3.

Mean of population "(\mu)" = "\dfrac{7+10+14+17+20}{5}=\dfrac{68}{5}"

The number of possible samples which can be drawn without replacement is "^{N}C_n=^{5}C_3=10."

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 no & Sample & Sample \\
& & mean\ (\bar{x})
\\ \hline
 1 & 7,10,14 & 31/3\\
 \hdashline
 2 & 7,10,17 & 34/3 \\
 \hdashline
 3 & 7,10,20 & 37/3 \\
 \hdashline
 4 & 7,14,17 & 38/3 \\
 \hdashline
 5 & 7,14,20 & 41/3 \\
 \hdashline
 6 & 7,17,20 & 44/3 \\
 \hdashline
 7 & 10,14,17 & 41/3 \\
 \hdashline
 8 & 10,14,20 & 44/3 \\
 \hdashline
 9 & 10,17,20 & 47/3 \\
 \hdashline
 10 &14,17,20 & 51/3 \\
 \hdashline
\end{array}"




"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f(\bar{X}) &\bar{X} f(\bar{X}) & \bar{X}^2f(\bar{X})
\\ \hline
 31/3 & 1/10 & 31/30 & 961/90 \\
 \hdashline
 34/3 & 1/10 & 34/30 & 1156/90 \\
 \hdashline
 37/3 & 1/10 & 37/30 & 1369/90 \\
 \hdashline
 38/3 & 1/10 & 38/30 & 1444/90 \\
 \hdashline
 41/3 & 2/10 & 82/30 & 3362/90 \\
 \hdashline
 44/3 & 2/10 & 88/30 & 3872/90 \\
 \hdashline
 47/3 & 1/10 & 47/30 & 2209/90 \\
 \hdashline
 51/3 & 1/10 & 51/30 & 2601/90 \\
 \hdashline
\end{array}"


Mean of sampling distribution 

"\mu_{\bar{X}}=E(\bar{X})=\sum\bar{X}_if(\bar{X}_i)=\dfrac{408}{30}=\dfrac{68}{5}=\mu"